Accountancy

Rajesh withdrawn rs12000 regularly the middle of each month. Interest is 12% per annum calculate interest on drawings 
Answer #1

Interest on drawings = 12000×12=144000

144000×6/12×12/100

1440×1/2×12=8640

answered by: Advika